# A Clique Partitioning-Based Algorithm for Graph Compression

### v2

@@ 5 Experimental Results
Speedup relative to FM. We define the speedup relative to FM as the ratio of the runing time of FM over CPGC. Figures 4d-f illustrate the speedup achieved by CPGC for bipartite graphs with $n$ = 32, 64, and 128 vertices in each partition, $m$ = 819 to 16 thousand for densities of 0.80, 0.90, and 0.98, and $\delta$ values ranging from 0.5 to 1. The running time of CPGC is consistently lower than that of FM, resulting in speedup by CPGC across all cases. This is primarily due to the selection of vertices in the while loop (Lines 4-6) in CSA, facilitating faster compression by extracting more than one i.e., $\gamma$, $\delta$-cliques in a single iteration, as shown in Figure 2(b). The speedup increases notably with higher $\delta$ and density. For example, with $n=32$ and $m=819$ for density 0.8, CPGC achieves an average speedup of 3.66 for $\delta=0.5$, increasing to 16.28 for $\delta=1$. Similarly, with $n=32$ and $\delta=1$, CPGC achieves an average speedup ranging from 16.28 for $m=819$ with density 0.8 to 20.97 for $m=16$ thousand with density 0.98. It is worth noting that for small graphs, certain combinations of $n$, $m$, and $\delta$ yield higher speedup values. For instance, with $n=128$ and $m=819$ for density 0.8, the speedup increases from 52.81 to 81.97 for $\delta=0.8$, and then slightly decreases to 60.68 for $\delta=1$.
5.3 Speeding-up Matching Algorithms.
Since the compressed representation of the original graph maintains all the path information, it can be used as input to other graph algorithms such as maximum cardinality matching, to speedup their execution. Consider Dinitz’s algorithm [6] for finding the maximum cardinality matching in a bipartite graph, whose running time is in $O(\sqrt{n}m)$. According to Theorem 4.3 the compression ratio of CPGC is $k=\Omega\Big{(}\frac{\delta\log n}{\log(2n^{2}/{m})}\Big{)}$, where $10<\delta<$ 1. This is the same as the compression ratio achieved by FM [9]. Because the number of edges in the compressed graph is reduced by a factor of $k$, using CPGC as preprocessing for Dinitz’s algorithm leads to a total time of $O(\sqrt{n}m/k+mn^{\delta})$. For $\delta<1/2$, the total running time becomes $O(\sqrt{n}m\log_{n}{(\frac{n^{2}}{m}}))$. To the best of our knowledge, this matches the best known asymptotic bound for maximum cardinality bipartite matching algorithms (also achieved by employing FM instead of CPGC and by the bipartite matching algorithm based on push-relabel method [10]). Similar reasoning applies if instead of Dinitz’s algorithm we consider the Hopcroft-Karp algorithm [11].
Figure 5 shows the average running time of Dinitz’s algorithm on the original bipartite graph $G$ with $n=32$ thousand vertices in each bipartition, and $m$ ranging from 214.75 million to 1.05 billion edges for densities = $0.80,0.90,$ and $0.98$ and different $\delta$. It also shows the run time of Dinitz’s algorithm executed on the compressed tripartite graph $G^{*}$, for different values of $\delta$, and densities $=0.80,0.90,$ and $0.98$. We observe that in all cases the Dinitz algorithm applied to the compressed graph $G^{*}$ outperforms its execution on the original graph $G$, achieving a reduction in the running time of up to 72.83%. Even when considering the preprocessing time required to compress the graph $G$, the running time of the Dinitz algorithm for large, dense graphs is still reduced by up to 25.09% with the compressed graph $G^{*}$. This reduction in the running time is directly proportional to the level of compression achieved, where larger compression ratios result in reduced running time. For instance, as discussed in the context of large graphs (Figure 3(a-c)), for a density of $0.98$, CPGC demonstrates an increasing compression ratio as the edge count $m$ reaches 1.05 billion with increasing values of $\delta$. Consequently, the run time of Dinitz’s algorithm on the compressed graph $G^{*}$ also decreases.
@@ 6 Conclusion and Future Work
We proposed a new algorithm for graph compression based on partitioning the graph into bipartite cliques. The proposed algorithm runs in time $O(mn^{\delta})$ which improves over the $O(mn^{\delta}\log^{2}n)$ time of the Feder-Motwani (FM) algorithm (the best existing clique partitioning-based graph compression algorithm). Our algorithm achieves an average compression ratio at least as good as that of FM. Experimental results showed that our algorithm achieved a compression ratio of up to 26% greater than that obtained by FM and executes up to 105.18 times faster than FM. We also investigated the reduction in running time of a cardinality matching algorithm (Dinitz’s algorithm) when using the compressed graph instead of the original graph. The results showed that for sufficient large dense graphs, using the compressed graph obtained by CPGC as an input to the matching algorithm leads to a reduction in the running time of up to 72.83% over the matching algorithm using the original graph. In future work, we plan to continue investigating the graph compression problem and propose algorithms that improve over the compression ratio obtained by our proposed algorithm.
A Appendix
A.1 FM algorithm.
In Algorithm 3 we give a high-level description of the FM algorithm. The FM algorithm first constructs neighborhood trees for each vertex $u_{i}$ in $U$, which are labeled binary trees of depth $r$, whose nodes are labeled by a bit string $\omega$, where $0\leq|\omega|\leq r$. The root node of a neighborhood tree contains all the vertices in $W$ and is labeled by the empty string, $\epsilon$. The bit string $\omega$ of a node is obtained by starting at the root and following the path in the tree up to the node and concatenating a 0 or a 1 to $\omega$ each time the path visits the left, or the right child, respectively. The nodes at level $i$ in the tree correspond to a partition of $W$ into sets of size $|W|/2^{i}$. Thus, each leaf node corresponds to a vertex in $W$. Each node in the neighborhood tree of $u_{i}$ stores $d_{i,\omega}$, the number of neighbors of $u_{i}$ in the set $W_{\omega}$ of vertices of $W$ that are associated with the node with label $\omega$, i.e., $d_{i,\omega}=|N(u_{i})\cap W_{\omega}|$, where $N(u_{i})$ is the set of the neighbors of $u_{i}$. In Figure 1(b), we show the neighborhood tree of vertex $u_{2}$ for the bipartite graph with $|U|=|W|=n=8$ given in Figure 1(a). Vertex $w_{3}$ is a leaf and would have $\omega=010$ in all neighborhood trees of the graph.
@@ A Appendix
6:     Starting with $t=1$ and $U_{t}=U$, iterate while $t\leq\hat{k}$ to select $\hat{k}$ vertices based on $c_{0}$ and $c_{1}$ (calculated as in Equation A.1) to form the ordered set $K_{q}=\{y_{1},y_{2},\ldots,y_{t}\}\subseteq W$. At each stage $t$, obtain set $U_{t+1}=\left\{u\in U_{t}\mid y_{1},\ldots,y_{t}\in N(u)\right\}$ which omits vertices from $U$ that are not adjacent to $K_{q}$. And update the corresponding neighborhood trees such that $d_{i,\omega}$ of the nodes along the path from root to the selected vertex is decreased by 1 for each neighborhood tree corresponding to $U_{t}$.
@@
10:For each clique in $\mathcal{C}=\{C_{1},\ldots,C_{q}\}$ add edges from all the vertices in the left partition $U_{q}$ to an additional vertex $z_{q}$, and from $z_{q}$ to all the vertices in the right partition $K_{q}$. This would compress the graph by replacing $|U_{q}|\times|K_{q}|$ edges with $|U_{q}|+|K_{q}|$ edges.
11:The remaining edges in $G$ are trivial cliques and are added in the compressed graph $G^{*}$.
12:Output: $G^{*}$, the compressed graph of $G$.
The algorithm computes $\hat{k}$ and then, in Line 5, it performs $\hat{k}$ iterations (indexed by $t$) to select vertices for the right partition of clique $C_{q}$. In each iteration it selects vertices at leaf nodes of the neighborhood trees by following a path based on $d_{i,\omega}$ of child nodes and the number of distinct ordered subsets of the neighbourhood tree at each level. In order to select a path at each level, the algorithm calculates $c_{0}$ and $c_{1}$ as follows:
(A.1) $c_{j}=C\left(U_{t},\mathscr{C}_{t,\omega\cdot j}\right)=\sum_{i:{u_{i}\in U_{t}}}d_{i,\omega\cdot j}\cdot\left(d_{i}-1\right)^{[{k}-t]},\,j=0,1.$
@@
The FM algorithm compresses the graph by adding a new vertex set $Z$ to the bipartite graph, thus converting it into a tripartite graph (Figure 1(c)), where each vertex in $Z$ corresponds to one clique extracted by the algorithm. Each vertex of the left and right partitions of $\delta$-clique $C_{q}$ is then connected via an edge to a new vertex $z_{q}\in Z$, thus forming a tripartite graph. This decreases the number of edges from $|U_{q}|\times|K_{q}|$ to $|U_{q}|+|K_{q}|$, thus compressing the graph. The compressed graph obtained by the algorithm preserves the path information of the original graph. The running time of the FM algorithm is $O(mn^{\delta}\log^{2}n)$. The FM algorithm can be extended to the case of non-bipartite graphs, where it computes a compression of a graph in time $O(mn^{\delta}\log^{2}n)$, as shown by Feder and Motwani [9].
A.2 Example: CPGC Execution.
We now show how CPGC works on a bipartite graph $G$ with partitions $U$ and $W$, $n=|U|=|W|=8$, and $|E|=54$, shown in Figure 6(a). CPGC (Algorithm 1), first initializes $q=0,n=|W|=|U|=8,\hat{m}=|E|=54,\mathcal{C}=\emptyset$, and $d_{w}=[0]_{n}$. It then calculates the degree of each vertex $d_{w_{j}}$ in Lines 2-4, as shown in Table 1. It calculates $\hat{k}(n,m,\delta)=2$ in Line 5. Thus, the condition for the while loop in Line 6 is met and CPGC calls CSA in Line 7. With the given inputs, CSA sorts $d_{w}$ in non-increasing order. Let $\{w_{4},w_{2},w_{3},w_{5},w_{6},w_{1},w_{7},w_{8}\}$ be the non-increasing order with corresponding degrees given as $d_{w_{\pi(j)}}$ in Table 1.
CPGC step $j$ 1 2 3 4 5 6 7 8

### v3

@@ 5.1 Results for Small Bipartite Graphs
Speedup relative to FM. We define the speedup relative to FM as the ratio of the runing time of FM over that of CPGR. Figures 3d-f illustrate the speedup achieved by CPGR for bipartite graphs with $n$ = 32, 64, and 128 vertices in each partition, $m$ = 819 to 16 thousand for densities of 0.80, 0.90, and 0.98, and $\delta$ values ranging from 0.5 to 1. The running time of CPGR is consistently lower than that of FM, resulting in speedup by CPGR across all cases. This is primarily due to the selection of vertices in the while loop (Lines 4-6) in CSA, facilitating greater $R$ by extracting more than one (i.e., $\gamma$), $\delta$-cliques in a single iteration, as shown in Figure 2b. The speedup increases notably with higher $\delta$ and density. For example, with $n=32$ and $m=819$ for density 0.8, CPGR achieves an average speedup of 3.66 for $\delta=0.5$, increasing to 16.28 for $\delta=1$. Similarly, with $n=32$ and $\delta=1$, CPGR achieves an average speedup ranging from 16.28 for $m=819$ with density 0.8 to 20.97 for $m=16$ thousand with density 0.98. It is worth noting that for small graphs, certain combinations of $n$, $m$, and $\delta$ yield higher speedup values. For instance, with $n=128$ and $m=819$ for density 0.8, the speedup increases from 52.81 to 81.97 for $\delta=0.8$, and then slightly decreases to 60.68 for $\delta=1$.
5.2 Results for Large Bipartite Graphs
We present the results obtained from testing CPGR on large bipartite graphs, where each bipartition consists of $n=2^{i}$ vertices, with $i$ ranging from 11 to 15, and a number of edges $m$ ranging from approximately 3.36 million to 1.05 billion.
Figure 4: CPGR on Large Graphs: Average reduction in the number of edges $R^{avg}$ ((a), (b), and (c)) and average running time ((d), (e), and (f)) for graphs with 3.36 million to 1.05 billion edges.
Average reduction in the number of edges ($R^{avg}$). Figures 4a-4c illustrate $R^{avg}$ for CPGR, defined as the percentage of edges removed from the original graph to obtain the restructured graph, i.e., $R^{avg}=\frac{m-m^{*}}{m}\cdot 100$. We observe a consistent trend: for fixed $n$ and $\delta$, $R^{avg}$ improves with increasing density. For example, when $m=214.75$ million edges at a density of $0.80$ and $\delta=0.5$, $R^{avg}$ is $48.72\%$, whereas $R^{avg}$ rises to $59\%$ for $m=1.05$ billion edges at a density of $0.98$. This correlation arises because higher density increases the likelihood of finding a large set of common neighbors for the right partition of a $\delta$-clique. Identifying more common neighbors allows CPGR to remove more edges, thereby improving $R^{avg}$.
However, when the density is held constant, $R^{avg}$ exhibits a non-monotonic behavior depending on $\delta$ and $n$. Consider Figure 4a, for $m=53.69$ million edges at density $0.80$, $R^{avg}$ initially increases from $45.6\%$ at $\delta=0.5$ to a peak of $50\%$ at $\delta=0.6$. Further increasing $\delta$ to $1$, however, leads to a decrease in $R^{avg}$ to $26.5\%$. As explained before, this is because a larger $\delta$ increases the size of the right partition $K_{q}$ of a $\delta$-clique, which consequently lowers the possibility of finding common neighbors.
Average running time. For a fixed $\delta$ and density, the running time of CPGR increases with the number of vertices, consistent with its $O(mn^{\delta})$ complexity. Conversely, for a fixed $n$ and density, the running time decreases as $\delta$ increases. For example, at density $0.80$ ($m=214.75$ million), the runtime drops from $176.27$ s at $\delta=0.5$ to $117.8$ s at $\delta=1$. This is because a larger $\delta$ limits the search for common neighbors, leading to lower reduction in the number of edges at lower densities and allowing CPGR to skip more vertices. Consequently, the expected increase in runtime with $\delta$ is not observed. Notably, even at a high density of $0.98$ ($m=1.05$ billion), the running time still decreases with increasing $\delta$ (from $164.83$ s at $\delta=0.5$ to $150.97$ s at $\delta=1$). This is attributed to the extraction of larger $\delta$-cliques at higher densities and $\delta$, which reduces the overall number of iterations in CPGR, thus lowering the runtime.
5.3 Results for General Graphs
Figure 5: CPGR on General Graphs: (a) average reduction in number of edges $R^{avg}$ and (b) average running time for general graphs with 409 million to 716 million edges.
In this section, we analyze CPGR’s performance on synthetically generated general graphs with $n=32000$ vertices and number of edges $m$ ranging from approximately 409 million to 716 million as explained in Section 5. We observe that $R^{avg}$ and running time follow the same trend with varying density and $\delta$ as seen in the large bipartite graphs.
Average reduction in the number of edges ($R^{avg}$). Figure 5a shows that for a fixed $\delta$, $R^{avg}$ improves with density. For instance, at $\delta=0.6$, $R^{avg}$ increases from $12.3\%$ for $m\approx 409$ million at density $0.4$ to $28.6\%$ for $m\approx 716$ million at density $0.7$, consistent with the increased likelihood of finding common neighbors. Similar to large bipartite graphs, $R^{avg}$ is non-monotonic with $\delta$ at a constant density. For $m\approx 716$ million at density $0.7$, $R^{avg}$ peaks at $33.3\%$ for $\delta=0.6$, after initially increasing from $28.6\%$ at $\delta=0.5$, before dropping to $9.1\%$ at $\delta=1$.
Average running time. Running time exhibits a non-monotonic trend with increasing $\delta$ and density. Larger $\delta$ leads to larger right partitions, reducing the likelihood of finding common neighbors in low-density graphs, thus fewer extracted cliques with larger partitions. Conversely, low $\delta$ (e.g., 0.5) results in smaller right partitions, increasing the chance of finding common neighbors, leading to higher $R$ and longer runtime, as evident in Figure 5b. The highest runtime (59.96 s) corresponds to the highest $R^{avg}$ ($33.3\%$) at density $0.7$ and $\delta=0.6$, while the runtime reduces to 21.21 s when $\delta=1$, and $R^{avg}$ drops to $9.1\%$ at the same density.
Table 2: Real-world Datasets (SuiteSparse Matrix Collection [9])
Sr. No Dataset name Rows Columns Edges Density
1 beacxc 497 506 50409 0.2
2 beaflw 497 507 53403 0.2
3 cari 400 1200 152800 0.3
4 mbeacxc 496 496 49920 0.2
5 mbeaflw 496 496 49920 0.2
Figure 6: CPGR on Real-world graphs: (a) average reduction in the number of edges ($R^{avg}$) and (b) average running time.
5.4 Results for Real-world Graphs
In this section, we analyze CPGR performance on low-density (up to 0.3), symmetric and asymmetric real-world graphs shown in Table 2. Given that $\delta<0.8$ would yield $\hat{k}=1$ and thus no reduction in the number of edges for these sparse graphs, we tested CPGR with $\delta\in\{0.8,0.9,1\}$.
Average reduction in the number of edges ($R^{avg}$). Figure 6a demonstrates that CPGR achieves $R^{avg}$ ranging from $18.7\%$ to $46.8\%$ even for the relatively low-density real-world graphs examined. Notably, for dataset cari, $R^{avg}$ remains consistent across all tested $\delta$ values. This stability is due to low density, which limits the maximum value of $\hat{k}$ to 2, thereby resulting in same the $R^{avg}$ regardless of $\delta$. While for other real-world graphs, $R^{avg}$ increases for $\delta$ = 1, this is because, when $\delta=1$, $\hat{k}$ larger for longer iterations of CPGR (as shown in Figure 2b), resulting in increase in $R^{avg}$.
Average running time. The running time of CPGR on the real-world datasets is presented in Figure 6b. Similar to $R^{avg}$, the running time generally remains stable for most datasets, which is due to the same reason of low density limiting $\hat{k}$. However, we observe for the datasets that show increase in $R^{avg}$ at $\delta$ = 1, has increase in running time, For example, the average running time for dataset beaflw increases from 0.007 s at $\delta=0.8$ and $0.9$ to 0.014 s at $\delta=1$. This increase in running time is because at $\delta=1$, CPGR could extract larger $\delta$-cliques (with $\hat{k}=2$), leading to higher $R^{avg}$ and thus a longer running time.
5.5 Speeding-up Graph Algorithms
In this section, we show the speedup achieved by downstream graph algorithms when we use CPGR algorithm as a preprocessing step across different graph types. Specifically, we present the speedup of the APSP algorithm on real-world datasets and the speedup of Dinitz’s algorithm on large synthetic bipartite datasets. Across these algorithms and dataset types, utilizing graphs with reduced number of edges obtained by CPGR as input consistently results in smaller running times.
Figure 7: CPGR + APSP($G^{*}$) vs. APSP($G$): Average running time of APSP algorithm on the original graph $G$ (labeled APSP($G$)) and on the restructured graph $G^{*}$ (labeled APSP($G^{*}$)), for large graphs with approximately 32,000 vertices and 256 million to 360 million edges corresponding to densities = $0.50,0.60,$ and $0.70$ and different $\delta$.
Speeding-up All-Pairs-Shortest-Path Algorithms. We evaluate the effectiveness of CPGR for speeding up Breadth-First Search (BFS)–based All-Pairs Shortest Path (APSP) computations on unweighted general graphs. Figure 7 shows the running time of APSP on the original graph $G$ and on the restructured graph $G^{*}$ obtained via CPGR. For each density, the bars represents the average cost of CPGR, APSP($G^{*}$), and APSP($G$), measured over 10 runs. To validate correctness, we compare the number of connected vertices in the original and restructured graphs. This validation confirms that vertex-to-vertex reachability is preserved between $G$ and $G^{*}$. Moreover, any path discovered by BFS in the restructured graph $G^{*}$ can be mapped back to a corresponding path in the original graph $G$ by contracting the length-two paths introduced by CPGR for $\delta$-cliques. Therefore, while path lengths in $G^{*}$ may differ, the APSP results remain correct with respect to the original graph.
Figure 7 shows the results on synthetic graphs with $n=32000$ vertices, varying both graph density and the parameter $\delta$. The results show that, across all tested densities, the combined execution time of CPGR and APSP($G^{*}$) consistently outperforms APSP($G$) on large graphs. Although the preprocessing cost increases with density, the reduction in APSP runtime offsets this overhead, yielding speedups of up to 1.74x. These results demonstrate that CPGR remains effective even for large, moderately dense graphs.
Speeding-up Matching Algorithms.
Figure 8: CPGR + Dinitz($G^{*}$) vs. Dinitz($G$): Average running time of Dinitz’s algorithm on the original bipartite graph $G$ (labeled Dinitz($G$)) and on the restructured graph $G^{*}$ (labeled Dinitz($G^{*}$)), for a large graph with approximately 32,000 vertices in each bipartition and 819.75 million to 1.05 billion edges corresponding to densities = $0.80,0.90,$ and $0.98$ and different $\delta$.
The restructured graph retains path information, enabling its use as input for other graph algorithms like maximum cardinality matching to accelerate execution. For bipartite matching, we focus on Dinitz’s algorithm, which computes a maximum matching via the standard unit-capacity flow formulation. In this formulation, each edge has unit capacity and feasible flows correspond exactly to matching. Under the CPGR restructuring, each edge $(u_{i},w_{j})$ belonging to an extracted $\delta$-clique is replaced by a two-edge path $(u_{i},z_{q},w_{j})$ in the restructured graph $G^{*}$, where both edges have unit capacity. This transformation preserves the set of feasible flows: a flow can traverse $(u_{i},z_{q},w_{j})$ in $G^{*}$ if and only if the corresponding edge $(u_{i},w_{j})$ could be used in $G$, and matching constraints continue to be enforced by the unit-capacity edges incident to vertices in the left and right partitions. Consequently, Dinitz’s algorithm computes a maximum flow of the same value on $G^{*}$ as on $G$, and the resulting matching can be mapped back to a maximum matching in $G$ in linear time.
Dinitz’s algorithm [11], with a runtime of $O(\sqrt{n}m)$ for bipartite matching, benefits from CPGR preprocessing. Given CPGR’s reduction in the number of edges $k=\Omega\Big(\frac{\delta\log n}{\log(2n^{2}/{m})}\Big)$ (matching FM’s), the total time becomes $O(\sqrt{n}m/k+mn^{\delta})$, which simplifies to $O(\sqrt{n}m\log_{n}{(\frac{n^{2}}{m}}))$ for $\delta<1/2$. Notably, this matches the asymptotic bound for maximum cardinality bipartite matching, achieved by both FM and push-relabel methods. Similar speedups are expected with the Hopcroft-Karp algorithm [16]. Better asymptotic bounds (almost linear in $m$) are achieved by recent bipartite cardinality matching algorithms such as those presented in [8] ($n^{2+o(1)}$-time randomized algorithm) and [21] ($\tilde{O}(m+n^{1.5})$-time randomized algorithm). These algorithms are mainly of theoretical interest, providing the best bounds, and not of practical interest. This is due to the use of sophisticated techniques that require significant implementation effort and they might not even be feasible to implement in practice. Thus, for our experimental analysis we use Dinitz’s algorithm as the baseline for comparisons.
Figure 8 compares the running time of Dinitz’s algorithm on the original bipartite graph $G$ (n=32n=32k per partition, $m=214.75$ Million to 1.05 Billion edges, densities 0.80-0.98) with the combined time of CPGR preprocessing and Dinitz on the restructured graph $G^{*}$. In all cases, the combined time outperforms Dinitz on $G$, achieving up to a 51.63% reduction (2.07x speedup). On average CPGR + Dinitz ($G^{*}$) achieves 1.69x speedup when compared to Dinitz ($G$). The speedup correlates with the reduction in the number of edges; for example, at density 0.98, the increasing reduction in the number of edges by CPGR with higher $m$ and $\delta$ leads to a faster Dinitz on $G^{*}$, resulting in up to a 96.07% reduction in Dinitz’s runtime compared to $G$.
@@ 6 Conclusion and Future Work
We proposed a novel Clique Partition based Graph Restructuring (CPGR) algorithm that operates by partitioning the graph into bipartite cliques. The algorithm has a time complexity of $O(mn^{\delta})$. Experimental results demonstrate the effectiveness of CPGR, achieving a reduction in the number of edges by up to 21% greater and execution times up to 105.18 times faster than the baseline FM algorithm. Furthermore, we investigated the impact of using the CPGR-restructured graph as input for a cardinality matching algorithm (Dinitz’s algorithm) and an All-Pairs Shortest Path (APSP) algorithm. Experimental results show that for sufficiently large dense graphs, our approach leads to a reduction in the total running time (including CPGR time) of up to 51.63%, resulting in a speedup of 2.07x over the matching algorithm on the original graph. Similarly, for the APSP algorithm, we observed up to 34.78% reduction in total running time, achieving a speedup of up to 1.74x. These speedups for fundamental graph algorithms highlight the practical benefits of our graph restructuring technique, particularly for large-scale real-world applications like route planning and social network analysis where faster processing is critical for efficiency and responsiveness.
Looking ahead, our results indicate that the edge reduction achieved by CPGR is influenced by the interaction between the graph size, the graph density, and the parameter $\delta$. A deeper understanding of these interactions, including the behavior observed for certain dense graphs, remains an open problem. Future work will focus on theoretically characterizing these interactions, developing adaptive strategies for selecting $\delta$, and extending CPGR to dynamic and streaming graph settings.
Appendix A Appendix
[…]
Figure 9: Neighborhood tree of vertex $u_{2}\in U$ that shows the path $\omega$ taken from root to the vertex $w_{j}\in W$ at the leaf and number of edges $d_{u_{2},\omega}$ for each node using the tuple $(\omega,d_{u_{2},\omega})$.
A.1 FM algorithm
In Algorithm 3 we give a high-level description of the FM algorithm. The FM algorithm first constructs neighborhood trees (Figure 9) for each vertex $u_{i}$ in $U$, which are labeled binary trees of depth $r$, whose nodes are labeled by a bit string $\omega$, where $0\leq|\omega|\leq r$. The root node of a neighborhood tree contains all the vertices in $W$ and is labeled by the empty string, $\epsilon$. The bit string $\omega$ of a node is obtained by starting at the root and following the path in the tree up to the node and concatenating a 0 or a 1 to $\omega$ each time the path visits the left, or the right child, respectively. The nodes at level $i$ in the tree correspond to a partition of $W$ into sets of size $|W|/2^{i}$. Thus, each leaf node corresponds to a vertex in $W$. Each node in the neighborhood tree of $u_{i}$ stores $d_{i,\omega}$, the number of neighbors of $u_{i}$ in the set $W_{\omega}$ of vertices of $W$ that are associated with the node with label $\omega$, i.e., $d_{i,\omega}=|N(u_{i})\cap W_{\omega}|$, where $N(u_{i})$ is the set of the neighbors of $u_{i}$. In Figure 9, we show the neighborhood tree of vertex $u_{2}$ for the bipartite graph with $|U|=|W|=n=8$ given in Figure 2. Vertex $w_{3}$ is a leaf and would have $\omega=010$ in all neighborhood trees of the graph.
@@ A.1 FM algorithm
6:   Starting with $t=1$ and $U_{t}=U$, iterate while $t\leq\hat{k}$ to select $\hat{k}$ vertices based on $c_{0}$ and $c_{1}$ (calculated as in Equation 3) to form the ordered set $K_{q}=\{y_{1},y_{2},\ldots,y_{t}\}\subseteq W$. At each stage $t$, obtain set $U_{t+1}=\left\{u\in U_{t}\mid y_{1},\ldots,y_{t}\in N(u)\right\}$ which omits vertices from $U$ that are not adjacent to $K_{q}$. And update the corresponding neighborhood trees such that $d_{i,\omega}$ of the nodes along the path from root to the selected vertex is decreased by 1 for each neighborhood tree corresponding to $U_{t}$.
@@
10:For each clique in $\mathcal{C}=\{C_{1},\ldots,C_{q}\}$ add edges from all the vertices in the left partition $U_{q}$ to an additional vertex $z_{q}$, and from $z_{q}$ to all the vertices in the right partition $K_{q}$. This would restructure the graph by replacing $|U_{q}|\times|K_{q}|$ edges with $|U_{q}|+|K_{q}|$ edges.
11:The remaining edges in $G$ are trivial cliques and are added in the restructured graph $G^{*}$.
12:Output: $G^{*}$, the restructured graph of $G$.
The algorithm computes $\hat{k}$ and then, in Line 5, it performs $\hat{k}$ iterations (indexed by $t$) to select vertices for the right partition of clique $C_{q}$. In each iteration it selects vertices at leaf nodes of the neighborhood trees by following a path based on $d_{i,\omega}$ of child nodes and the number of distinct ordered subsets of the neighborhood tree at each level. In order to select a path at each level, the algorithm calculates $c_{0}$ and $c_{1}$ as follows:
$c_{j}=C\left(U_{t},\mathscr{C}_{t,\omega\cdot j}\right)=\sum_{i:{u_{i}\in U_{t}}}d_{i,\omega\cdot j}\cdot\left(d_{i}-1\right)^{[{k}-t]},\,j=0,1.$ (3)
@@
The FM algorithm restructures the graph by adding a new vertex set $Z$ to the bipartite graph, thus converting it into a tripartite graph (Figure 2a), where each vertex in $Z$ corresponds to one clique extracted by the algorithm. Each vertex of the left and right partitions of $\delta$-clique $C_{q}$ is then connected via an edge to a new vertex $z_{q}\in Z$, thus forming a tripartite graph. This decreases the number of edges from $|U_{q}|\times|K_{q}|$ to $|U_{q}|+|K_{q}|$, thus reducing the number of edges in the graph. The restructured graph obtained by the algorithm preserves the path information of the original graph. The running time of the FM algorithm is $O(mn^{\delta}\log^{2}n)$. The FM algorithm can be extended to the case of non-bipartite graphs, where it restructures a graph in time $O(mn^{\delta}\log^{2}n)$, as shown by Feder and Motwani [14].
A.2 Example: CPGR Execution
We now show how CPGR works on a bipartite graph $G$ with partitions $U$ and $W$, $n=|U|=|W|=8$, and $|E|=54$, shown in Figure 10(a). CPGR (Algorithm 1), first initializes $q=0,n=|W|=|U|=8,\hat{m}=|E|=54,\mathcal{C}=\emptyset$, and $d_{w}=[0]_{n}$. It then calculates the degree of each vertex $d_{w_{j}}$ in Lines 6-8, as shown in Table 3. It calculates $\hat{k}(n,m,\delta)=2$ in Line 9. Thus, the condition for the while loop in Line 10 is met and CPGR calls CSA in Line 11. With the given inputs, CSA sorts $d_{w}$ in non-increasing order. Let $\{w_{4},w_{2},w_{3},w_{5},w_{6},w_{1},w_{7},w_{8}\}$ be the non-increasing order with corresponding degrees given as $d_{w_{\pi(j)}}$ in Table 3.
CPGR step $j$ 1 2 3 4 5 6 7 8